A man buys a car for $34,000. If the interest rate on the loan is 12%, compounded monthly, and if he wants to make monthly payments of $900 for 36 months, how much must he put down?
Answer #1

Down payment = Cost of Car - PV of car payments

PV of car payments can be calculated using formula:

PV =

[1-(4+5)-) [1-(1+r)-n LT P= Periodic Payment r=rate per period n = number of periods

r = 12%/12 = 1% (monthly)

n = 36

PV = 900 * 1-(1+0.01) - 36 0,01

PV = 900 * 30.1075

PV = $27,096.75

Down payment = $34,000 - $27,096.75

Down payment = $6,903.25